OpenStudy (anonymous):

Which is a solution to the equation y = 2x – 5? A. (3, 11) B. (2, 1) C. (1, –3) D. (5, –5)

OpenStudy (anonymous):

vague question

OpenStudy (anonymous):

thats exactly how they gave it to me :/

OpenStudy (anonymous):

ohhh i think i get it

OpenStudy (anonymous):

a,b,c,and d are numbers that could replace x and y so (x,y)

OpenStudy (anonymous):

right?

OpenStudy (youngstudier):

@jacob1728 what do you think the answer is?

OpenStudy (anonymous):

d?

OpenStudy (youngstudier):

Nope. Plug the answer choices into the equation. So, let's start with plugging in answer choice A. The equation is: y = 2x - 5 Plug in answer choice A And it's: 11 = 2*3 - 5 Now solve the right side of the equation. Does it equal 11?

OpenStudy (anonymous):

no

OpenStudy (youngstudier):

Yep, so let's plug in answer choice B to see if that works. 1 = 2*2 - 5 1 = ?

OpenStudy (anonymous):

nah its b because 2*2 is 4 and 4-5 is -1 i think

OpenStudy (anonymous):

its not*

OpenStudy (youngstudier):

Exactly! :D Next step is to plug in C. -3 = 2*1 - 5 -3 = ?

OpenStudy (anonymous):

sorry i accidentaly closed openstudy

OpenStudy (anonymous):

anyways..

OpenStudy (anonymous):

2*1=2-5 is -3 yah c is it

OpenStudy (youngstudier):

Correct! ^_^

OpenStudy (anonymous):

k can you help me w/ 4 more?(maybe less if i understand the questions)

OpenStudy (youngstudier):

Sure! ^-^ I'd be glad to.

OpenStudy (anonymous):

thanks!:D